OpenStudy (anonymous):

I really need the help Darian is buying an entertainment center for his living room wall. The tag on his first choice shows that the furniture is forty inches wide and ninety inches high. Darian has a space for it that is 5 feet wide and 7 feet high. Will his first choice fit? Explain.

OpenStudy (texaschic101):

12 inches = 1 ft. so, 40 in = 40/12 = 3 1/3 ft wide. 90 inches high. = 90/12 = 7 1/2 feet high His space that holds a t.v. is 5 foot wide and 7 foot high. This t.v. will not fit. The width will fit, but the height will not work. The t.v. is 6 inches too high.